10,30,15,16,48,24,25 what number is next

Answer by JulietG(1812)

10,30,15,16,48,24,25
3,2,1
Multiply by 3
Divide by 2
Add 1
10 * 3 = 30
30 / 2 = 15
15 + 1 = 16
16 * 3 = 48
48 / 2 = 24
24 + 1 = 25
25 * 3 = 75 <<<<------
75 / 2 = 37.5
37.5 + 1 = 38.5